At 5.5 percent interest, how long does it take to double your money? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

  Length of time years

At 5.5 percent interest, how long does it take to quadruple it? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  Length of time years
Answer #1

1)

FV=PV*(1+r)^n

2=1*(1+5.5%)^n

(1.055)^n=2

take natural log on both sides

n=ln(2)/ln(1.055)

n=12.95 years

the above is the answer

2)

FV=PV*(1+r)^n

4=1*(1+5.5%)^n

(1.055)^n=4

take natural log on both sides

n=ln(4)/ln(1.055)

n=25.89 years

the above is the answer